# The Orbits of Isolated Dwarfs in the Local Group from New 3D Kinematics: Constraints on First Infall, Backsplash, and Quenching Mechanisms

### v1

@@ 3.4 Orbital History Uncertainties
To understand how statistically common the orbital histories illustrated in Figure  2 are and to identify any backsplash galaxies, we must account for measurement uncertainties on the LOS velocity, distance, and PM in the numerical orbit integrations [^5] In this work, we do not test alternative MW and M31 masses and their effect on the orbits of the six dwarfs, as in Patel et al. (2020). However, as noted in our previous work, the assumed masses of the MW and M31 do significantly contribute to the orbital uncertainties. For consistency with B24, we only test a fixed set of masses in this work and leave further investigation of how different MW and M31 masses affect the result presented here to future work.
@@
From here on, first passage refers to galaxies that completed a pericenter recently about a more massive galaxy and are now moving away from the host galaxy (e.g., Leo I Boylan-Kolchin et al., 2013; Sohn et al., 2013), while first infall refers to galaxies where a pericenter has not yet occurred, but the galaxy is still moving towards a more massive galaxy. Backsplash is defined as one (or more) pericenters within the virial radius of the MW or M31 in the last 6 Gyr.
@@
For certain dwarfs, such as Aquarius, extrapolating from the orbits shown in the top row of Figure 3, it is possible that it interacted with the MW beyond 6 Gyr ago, but again, this is only speculation as we do not advise integrating beyond the 6 Gyr time period (see Bennet et al., 2024; Patel et al., 2020; Santistevan et al., 2020).
@@
Note. — Halo masses estimated from the Moster et al. (2013) abundance matching relation and Plummer scale lengths ($r_{P}$) derived by fitting the dynamical mass for each galaxy to a Plummer profile. Dynamical masses are computed with the Wolf et al. (2010) mass estimator. Stellar masses, half-light radii, and velocity dispersions are taken from McConnachie (2012). A Plummer scale length of 0.0001 kpc represents galaxies modeled as point masses.
@@
This Work This Work Buck et al. (2019)
@@
Note. — Probabilities of dwarf galaxies in our sample being backsplash galaxies compared to those computed using radial velocity and position in Buck et al. (2019) where available.
@@ 3.5 Comparison to Previous Orbits
Figure 4: Comparison of the absolute PM measurements for Leo T and Pegasus, the only two dwarfs in our sample with existing literature PMs. We show McConnachie & Venn (2020b) in blue diamonds, Battaglia et al. (2022a) in red circles, B24 in green triangles and our own results as black squares. As can be seen the PMs reported in this work are consistent with, though far more precise than, previous estimates. The different values reported using Gaia DR3 data are primarily the result of the different membership selection methods used by those works.
@@
For both objects, we report substantial improvements in the uncertainty here. This is partly because the time baseline is substantially larger, increasing from 1.84 yrs to 8.01 yrs for Pegasus, and 4.37 yrs to 11.03 yrs for Leo T. However, the greatest improvement is in the number of sources considered. In B24, the absolute reference frame is set by the stars in the Gaia catalog; however, in this work, the reference frame is set by background galaxies. For these dwarfs, there are substantially more background galaxies in the HST imaging than there are stars in the Gaia catalog, yielding a more certain anchor to the absolute reference frame. Finally, a larger number of sources also allows a more robust selection of member stars compared to background objects, lowering potential contamination and improving the measured uncertainties. This improvement also applies to previous Gaia measurements of Leo T and Pegasus’ PMs (McConnachie & Venn, 2020b; Battaglia et al., 2022a) for similar reasons. Figure 4 clearly illustrates the past PM measurements compared to our much improved PM measurements.
@@ 4 Discussion
In the following section, we examine our dwarf sample galaxy by galaxy, making specific comparisons to previous literature regarding their natures and evolutionary histories. For illustrative purposes, Figure 5 shows the literature SFHs for Aquarius (Cole et al., 2014), Cetus (Monelli et al., 2010a), and Leo T (Weisz et al., 2014).
Figure 5: Star formation histories for the past 10 Gyr for an illustrative subsample of our dwarfs. The y-axis shows the star formation rate in arbitrary units. Uncertainties are shown in the gray lines. Left panel: Aquarius from Cole et al. (2014), Aquarius shows relatively constant star formation between 2.5 and 7 Gyr ago, with a major burst at just beyond $\sim$7 Gyr ago. Center panel: Cetus from Monelli et al. (2010a), Cetus quenches at $\sim$8 Gyrs, the slight star formation around $\sim$4 Gyrs is likely caused by blue straggler stars. Right panel: Leo T from Weisz et al. (2012), Leo T shows relative constant star formation similar to Aquarius, but significantly more stochastic.
@@ 4.1 Aquarius
Aquarius is a very isolated dwarf at the edge of the LG, more than 1 Mpc from the MW or M31. It shows an SFH with a relatively constant star formation (see Figure 5) between 2.5 and 7 Gyr ago, and a slight decline in the star formation rate in recent times (this may be a result of increased time resolution in the SFH toward the present). Its most notable SFH feature is a large star formation burst at $\sim$7 Gyrs ago (Cole et al., 2014). Aside from this burst, which we will discuss below, this SFH is consistent with those of isolated dwarf galaxies in Aquarius’ mass range, as found in simulations (Garrison-Kimmel et al., 2019).
@@
Aquarius also has a large neutral hydrogen reservoir with a gas mass of $M_{HI}=4.1\pm 0.3\times 10^{6}\,\rm M_{\odot}$, (Putman et al., 2021) giving it a gas/stellar mass ratio of $M_{HI}/M_{*}\sim 1.5$, typical for an isolated dwarf galaxy (Huang et al., 2012).
@@
Despite its isolation at the edge of the LG, the burst of star formation $\sim$7 Gyrs ago may be due to an interaction with a more massive host, but our orbits exclude the possibility of any close interactions between Aquarius and the MW or M31. This starburst also coincides with a slight dip in the average metallicity of stars produced in this epoch compared to earlier stars, as well as a larger metallicity spread for stars older than 8 Gyrs (Cole et al., 2014). One possibility inferred from the combination of observed phenomena around $\sim$7 Gyrs ago is a dwarf-dwarf merger, such that Aquarius merged with another, smaller dwarf galaxy around this time.
@@
In B24, we discussed evidence of dwarf-dwarf mergers in Leo A ($\sim$6 Gyrs ago) and Leo I ($\sim$5 Gyrs ago) due to similar peaks in their SFHs that cannot be explained by interactions with the MW or M31. Similar star formation bursts pointing to possible merger events are also seen in Fornax dSph (del Pino et al., 2015; Rusakov et al., 2021) and Andromeda II (del Pino et al., 2017). However, in these cases, it is harder to separate the influence of the massive host (the MW or M31) from possible merger signals. Fornax dSph has been a long-term member of the MW satellite system (Fritz et al., 2018b; Rusakov et al., 2021; Battaglia et al., 2022b), and therefore has a more complicated interaction history with the MW that cannot be ignored. There are currently no published PMs for Andromeda II that would allow the kind of kinematic modeling needed to eliminate the possibility of the effects being caused by M31 or M33. If true, our results increase the number of possible dwarf-dwarf mergers in the LG at intermediate times to three instances.
@@
Other explanations for the star formation peaks include the inflow of pristine gas from the cosmic web or interactions with large-scale structure filaments, which could also trigger star formation bursts (Ledinauskas & Zubovas, 2018). Purely internal dynamical processes could also explain the peaks (Cole et al., 2007). While the metallicity features in these three dwarfs seem to favor a dwarf-dwarf merger hypothesis (Cole et al., 2014; Kirby et al., 2017; Ruiz-Lara et al., 2021), more data is needed to draw firmer conclusions.
@@
Simulations have shown that dwarf-dwarf mergers at intermediate times (4-8 Gyr)—such as our three possible mergers—occur from dwarf pairs that form and become bound together at early times (10-12 Gyrs) (Chamberlain et al., 2024a, b). Including our sample, there are 15 LG dwarfs in the literature with kinematics sufficiently precise to distinguish between a dwarf-dwarf merger and the influence of a massive host galaxy (e.g. Sohn et al., 2013, 2017, 2020; McConnachie et al., 2021; Battaglia et al., 2022b; Bennet et al., 2024). This implies a potential merger fraction of $\sim$20%, which matches that found in Chamberlain et al. (2024a, b) for low mass dwarfs. However, the mass range of the LG dwarfs examined in Chamberlain et al. (2024a, b) is lower, and the LG is a more crowded environment than those probed in those studies.
@@
We also find that compared to the $\sim 30$ classical dwarfs in the modern LG ($M_{V}\leq-8$ mag), we would expect to find $\sim 3$ pairs at z=0 based on the above work and a $\sim$20% merger fraction. If we compare this to the actual LG we find that there are fewer pairs than expected, with the LMC and SMC being the only kinematically confirmed pair in this mass range [^6] Other possible kinematic associations of dwarf galaxies in the MW halo such as the Leo-Crater group (Júlio et al., 2024) and the pair Pegasus III and Pisces II (Richstein et al., 2022) have been reported using 3D kinematics. However, these galaxies are ultra-faint dwarfs rather than the classical mass dwarfs as those explored in this work.. More work is needed to understand the frequency of dwarf-dwarf mergers across a wider range of dwarf masses in the LG.
@@ 4.2 Cetus
Cetus and Tucana have been used extensively as examples of so-called backsplash galaxies, i.e., galaxies that are outside the virial radius of a more massive galaxy or cluster at present, but retain evidence of a previous interaction (e.g. Teyssier et al., 2012; Taibi et al., 2018; Blaña et al., 2020; Santos-Santos et al., 2023).
More generally, this definition, in part, is also motivated by a lack of significant amounts of HI gas or recent star formation (see Figure 5). Cetus’ SFH shows substantial star formation after reionization, however, this stops around $\sim 8$ Gyrs ago as Cetus undergoes rapid quenching (Monelli et al., 2010a). Cetus and Tucana are the only isolated LG dwarfs that exhibit these properties, separating them from the rest of the isolated LG dwarfs, which are either transitional or irregular dwarfs, with evidence of recent star formation and HI reservoirs (Grcevich & Putman, 2009; Spekkens et al., 2014; Putman et al., 2021).
Furthermore, our GBT observations of Cetus (see Section 2.2) were a non-detection pushing the HI upper limit for this galaxy down to $8\times 10^{4}\,\rm M{{}_{\odot}}$ from the previous upper limit reported in Putman et al. (2021) ($9.6\times 10^{4}\,\rm M{{}_{\odot}}$). Therefore, we conclude that Cetus is a gas-poor dwarf with an HI upper limit of $\sim$3% of the reported stellar mass (McConnachie, 2012). This is the most robust HI gas limit in the LG outside the MW halo.
Though many studies have labeled Cetus as a backsplash galaxy based on these properties, some previous kinematic studies have indicated that these galaxies are more likely to be truly isolated (Fraternali et al., 2009; Buck et al., 2019). These works examine the radial velocity and distance of Cetus and Tucana from the M31 and the MW, and compare this to simulated dwarfs in LG analogs to explore the probability that a dwarf with the observed properties is either backsplash or on first infall. These show that dwarfs with Tucana and Cetus’ properties are likely (but not certain) to be on first infall.
@@
The right most panel of Figure 6 shows the distribution of $\rm r_{peri}$, the distance at pericenter, relative to the MW and M31, corresponding to the orbits in the left and middle panels. These distributions indicate that passages around the MW were more likely to occur at large distances (i.e., two times the MW’s virial radius, or $\approx$500 kpc). In contrast, there is a bimodal distribution in the $\rm r_{peri}$ distribution relative to M31 such that the first distribution peaks within M31’s virial radius of 329 kpc. Therefore, while the likelihood of a passage around M31 is less statistically common (44% of orbits), a close passage around M31 is more likely. Additionally, the burst of star formation around 8 Gyr ago is more consistent with the timing of a passage around M31 ($1\sigma$ for $\rm t_{peri}=5.32-8.94$ Gyr ago. Together, our orbital results combined with the Cetus SFH paint a consistent picture that it could be a backsplash galaxy of M31. Our conclusions are consistent with those of simulations (Garrison-Kimmel et al., 2019; Joshi et al., 2021; Samuel et al., 2022) showing that a dwarf with the mass of Cetus should not self-quench without the influence of a massive host (Garrison-Kimmel et al., 2019); though it readily quenches if such a host is present.
@@
A final possibility for Cetus if its quenching is not a result of an interaction with the MW or M31 is that it was once part of a group infall into the LG with the LMC, M33, or another LG galaxy that came in earlier than 6 Gyr ago, and was then tidally removed from its companion(s), leaving Cetus behind. Interactions during this scenario may plausibly explain Cetus’ SFH. Another consideration is that halo deformations due to the passage of massive satellites like the LMC and subsequent center of mass responses from massive host galaxies are not accounted for with the rigid potentials assumed in this work (see Garavito-Camargo et al., 2021; Patel et al., 2025). This phenomenon is beyond the scope of this analysis, but it is important to acknowledge that changing halo shape can alter orbital histories.

### v2

@@ 3.4 Orbital History Uncertainties
To understand how statistically common the orbital histories illustrated in Figure  2 are and to identify any backsplash galaxies, we must account for measurement uncertainties on the LOS velocity, distance, and PM in the numerical orbit integrations [^5] In this work, we do not test alternative MW and M31 masses and their effect on the orbits of the six dwarfs, as in E. Patel et al. (2020). However, as noted in our previous work, the assumed masses of the MW and M31 do significantly contribute to the orbital uncertainties. For consistency with B24, we only test a fixed set of masses in this work and leave further investigation of how different MW and M31 masses affect the result presented here to future work.
@@
From here on, first passage refers to galaxies that completed a pericenter recently about a more massive galaxy and are now moving away from the host galaxy (e.g., Leo I M. Boylan-Kolchin et al., 2013; S. T. Sohn et al., 2013), while first infall refers to galaxies where a pericenter has not yet occurred, but the galaxy is still moving towards a more massive galaxy. Backsplash is defined as one (or more) pericenters within the virial radius of the MW or M31 in the last 6 Gyr.
@@
For certain dwarfs, such as Aquarius, extrapolating from the orbits shown in the top row of Figure 3, it is possible that it interacted with the MW beyond 6 Gyr ago, but again, this is only speculation as we do not advise integrating beyond the 6 Gyr time period (see P. Bennet et al., 2024; E. Patel et al., 2020; I. B. Santistevan et al., 2020).
@@
Note. — Halo masses estimated from the B. P. Moster et al. (2013) abundance matching relation and Plummer scale lengths ($r_{P}$) derived by fitting the dynamical mass for each galaxy to a Plummer profile. Dynamical masses are computed with the J. Wolf et al. (2010) mass estimator. Stellar masses, half-light radii, and velocity dispersions are taken from A. W. McConnachie (2012). A Plummer scale length of 0.0001 kpc represents galaxies modeled as point masses.
@@
This Work This Work T. Buck et al. (2019)
@@
Note. — Probabilities of dwarf galaxies in our sample being backsplash galaxies compared to those computed using radial velocity and position in T. Buck et al. (2019) where available.
@@ 3.5 Comparison to Previous Orbits
Figure 4: Comparison of the absolute PM measurements for Leo T and Pegasus, the only two dwarfs in our sample with existing literature PMs. We show A. W. McConnachie & K. A. Venn (2020b) in blue diamonds, G. Battaglia et al. (2022a) in red circles, B24 in green triangles and our own results as black squares. As can be seen the PMs reported in this work are consistent with, though far more precise than, previous estimates. The different values reported using Gaia DR3 data are primarily the result of the different membership selection methods used by those works.
@@
For both objects, we report substantial improvements in the uncertainty here. This is partly because the time baseline is substantially larger, increasing from 1.84 yrs to 8.01 yrs for Pegasus, and 4.37 yrs to 11.03 yrs for Leo T. However, the greatest improvement is in the number of sources considered. In B24, the absolute reference frame is set by the stars in the Gaia catalog; however, in this work, the reference frame is set by background galaxies. For these dwarfs, there are substantially more background galaxies in the HST imaging than there are stars in the Gaia catalog, yielding a more certain anchor to the absolute reference frame. Finally, a larger number of sources also allows a more robust selection of member stars compared to background objects, lowering potential contamination and improving the measured uncertainties. This improvement also applies to previous Gaia measurements of Leo T and Pegasus’ PMs (A. W. McConnachie & K. A. Venn, 2020b; G. Battaglia et al., 2022a) for similar reasons. Figure 4 clearly illustrates the past PM measurements compared to our much improved PM measurements.
@@ 4 Discussion
In the following section, we examine our dwarf sample galaxy by galaxy, making specific comparisons to previous literature regarding their natures and evolutionary histories. For illustrative purposes, Figure 5 shows the literature SFHs for Aquarius (A. A. Cole et al., 2014), Cetus (M. Monelli et al., 2010a), and Leo T (D. R. Weisz et al., 2014).
Figure 5: Star formation histories for the past 10 Gyr for an illustrative subsample of our dwarfs. The y-axis shows the star formation rate in arbitrary units. Uncertainties are shown in the gray lines. Left panel: Aquarius from A. A. Cole et al. (2014), Aquarius shows relatively constant star formation between 2.5 and 7 Gyr ago, with a major burst at just beyond $\sim$7 Gyr ago. Center panel: Cetus from M. Monelli et al. (2010a), Cetus quenches at $\sim$8 Gyrs, the slight star formation around $\sim$4 Gyrs is likely caused by blue straggler stars. Right panel: Leo T from D. R. Weisz et al. (2012), Leo T shows relative constant star formation similar to Aquarius, but significantly more stochastic.
@@ 4.1 Aquarius
Aquarius is a very isolated dwarf at the edge of the LG, more than 1 Mpc from the MW or M31. It shows an SFH with a relatively constant star formation (see Figure 5) between 2.5 and 7 Gyr ago, and a slight decline in the star formation rate in recent times (this may be a result of increased time resolution in the SFH toward the present). Its most notable SFH feature is a large star formation burst at $\sim$7 Gyrs ago (A. A. Cole et al., 2014). Aside from this burst, which we will discuss below, this SFH is consistent with those of isolated dwarf galaxies in Aquarius’ mass range, as found in simulations (S. Garrison-Kimmel et al., 2019).
@@
Aquarius also has a large neutral hydrogen reservoir with a gas mass of $M_{HI}=4.1\pm 0.3\times 10^{6}\,\rm M_{\odot}$, (M. E. Putman et al., 2021) giving it a gas/stellar mass ratio of $M_{HI}/M_{*}\sim 1.5$, typical for an isolated dwarf galaxy (S. Huang et al., 2012).
@@
Despite its isolation at the edge of the LG, the burst of star formation $\sim$7 Gyrs ago may be due to an interaction with a more massive host, but our orbits exclude the possibility of any close interactions between Aquarius and the MW or M31. This starburst also coincides with a slight dip in the average metallicity of stars produced in this epoch compared to earlier stars, as well as a larger metallicity spread for stars older than 8 Gyrs (A. A. Cole et al., 2014). One possibility inferred from the combination of observed phenomena around $\sim$7 Gyrs ago is a dwarf-dwarf merger, such that Aquarius merged with another, smaller dwarf galaxy around this time.
@@
In B24, we discussed evidence of dwarf-dwarf mergers in Leo A ($\sim$6 Gyrs ago) and Leo I ($\sim$5 Gyrs ago) due to similar peaks in their SFHs that cannot be explained by interactions with the MW or M31. Similar star formation bursts pointing to possible merger events are also seen in Fornax dSph (A. del Pino et al., 2015; V. Rusakov et al., 2021) and Andromeda II (A. del Pino et al., 2017). However, in these cases, it is harder to separate the influence of the massive host (the MW or M31) from possible merger signals. Fornax dSph has been a long-term member of the MW satellite system (T. K. Fritz et al., 2018b; V. Rusakov et al., 2021; G. Battaglia et al., 2022b), and therefore has a more complicated interaction history with the MW that cannot be ignored. There are currently no published PMs for Andromeda II that would allow the kind of kinematic modeling needed to eliminate the possibility of the effects being caused by M31 or M33. If true, our results increase the number of possible dwarf-dwarf mergers in the LG at intermediate times to three instances.
@@
Other explanations for the star formation peaks include the inflow of pristine gas from the cosmic web or interactions with large-scale structure filaments, which could also trigger star formation bursts (E. Ledinauskas & K. Zubovas, 2018). Purely internal dynamical processes could also explain the peaks (A. A. Cole et al., 2007). While the metallicity features in these three dwarfs seem to favor a dwarf-dwarf merger hypothesis (A. A. Cole et al., 2014; E. N. Kirby et al., 2017; T. Ruiz-Lara et al., 2021), more data is needed to draw firmer conclusions.
@@
Simulations have shown that dwarf-dwarf mergers at intermediate times (4-8 Gyr)—such as our three possible mergers—occur from dwarf pairs that form and become bound together at early times (10-12 Gyrs) (K. Chamberlain et al., 2024a, b). Including our sample, there are 15 LG dwarfs in the literature with kinematics sufficiently precise to distinguish between a dwarf-dwarf merger and the influence of a massive host galaxy (e.g. S. T. Sohn et al., 2013, 2017, 2020; A. W. McConnachie et al., 2021; G. Battaglia et al., 2022b; P. Bennet et al., 2024). This implies a potential merger fraction of $\sim$20%, which matches that found in K. Chamberlain et al. (2024a, b) for low mass dwarfs. However, the mass range of the LG dwarfs examined in K. Chamberlain et al. (2024a, b) is lower, and the LG is a more crowded environment than those probed in those studies.
@@
We also find that compared to the $\sim 30$ classical dwarfs in the modern LG ($M_{V}\leq-8$ mag), we would expect to find $\sim 3$ pairs at z=0 based on the above work and a $\sim$20% merger fraction. If we compare this to the actual LG we find that there are fewer pairs than expected, with the LMC and SMC being the only kinematically confirmed pair in this mass range [^6] Other possible kinematic associations of dwarf galaxies in the MW halo such as the Leo-Crater group (M. P. Júlio et al., 2024) and the pair Pegasus III and Pisces II (H. Richstein et al., 2022) have been reported using 3D kinematics. However, these galaxies are ultra-faint dwarfs rather than the classical mass dwarfs as those explored in this work.. More work is needed to understand the frequency of dwarf-dwarf mergers across a wider range of dwarf masses in the LG.
@@ 4.2 Cetus
Cetus and Tucana have been used extensively as examples of so-called backsplash galaxies, i.e., galaxies that are outside the virial radius of a more massive galaxy or cluster at present, but retain evidence of a previous interaction (e.g. M. Teyssier et al., 2012; S. Taibi et al., 2018; M. Blaña et al., 2020; I. M. E. Santos-Santos et al., 2023).
More generally, this definition, in part, is also motivated by a lack of significant amounts of HI gas or recent star formation (see Figure 5). Cetus’ SFH shows substantial star formation after reionization, however, this stops around $\sim 8$ Gyrs ago as Cetus undergoes rapid quenching (M. Monelli et al., 2010a). Cetus and Tucana are the only isolated LG dwarfs that exhibit these properties, separating them from the rest of the isolated LG dwarfs, which are either transitional or irregular dwarfs, with evidence of recent star formation and HI reservoirs (J. Grcevich & M. E. Putman, 2009; K. Spekkens et al., 2014; M. E. Putman et al., 2021).
Furthermore, our GBT observations of Cetus (see Section 2.2) were a non-detection pushing the HI upper limit for this galaxy down to $8\times 10^{4}\,\rm M{{}_{\odot}}$ from the previous upper limit reported in M. E. Putman et al. (2021) ($9.6\times 10^{4}\,\rm M{{}_{\odot}}$). Therefore, we conclude that Cetus is a gas-poor dwarf with an HI upper limit of $\sim$3% of the reported stellar mass (A. W. McConnachie, 2012). This is the most robust HI gas limit in the LG outside the MW halo.
Though many studies have labeled Cetus as a backsplash galaxy based on these properties, some previous kinematic studies have indicated that these galaxies are more likely to be truly isolated (F. Fraternali et al., 2009; T. Buck et al., 2019). These works examine the radial velocity and distance of Cetus and Tucana from the M31 and the MW, and compare this to simulated dwarfs in LG analogs to explore the probability that a dwarf with the observed properties is either backsplash or on first infall. These show that dwarfs with Tucana and Cetus’ properties are likely (but not certain) to be on first infall.
@@
The right most panel of Figure 6 shows the distribution of $\rm r_{peri}$, the distance at pericenter, relative to the MW and M31, corresponding to the orbits in the left and middle panels. These distributions indicate that passages around the MW were more likely to occur at large distances (i.e., two times the MW’s virial radius, or $\approx$500 kpc). In contrast, there is a bimodal distribution in the $\rm r_{peri}$ distribution relative to M31 such that the first distribution peaks within M31’s virial radius of 329 kpc. Therefore, while the likelihood of a passage around M31 is less statistically common (44% of orbits), a close passage around M31 is more likely. Additionally, the burst of star formation around 8 Gyr ago is more consistent with the timing of a passage around M31 ($1\sigma$ for $\rm t_{peri}=5.32-8.94$ Gyr ago. Together, our orbital results combined with the Cetus SFH paint a consistent picture that it could be a backsplash galaxy of M31. Our conclusions are consistent with those of simulations (S. Garrison-Kimmel et al., 2019; G. D. Joshi et al., 2021; J. Samuel et al., 2022) showing that a dwarf with the mass of Cetus should not self-quench without the influence of a massive host (S. Garrison-Kimmel et al., 2019); though it readily quenches if such a host is present.
@@
A final possibility for Cetus if its quenching is not a result of an interaction with the MW or M31 is that it was once part of a group infall into the LG with the LMC, M33, or another LG galaxy that came in earlier than 6 Gyr ago, and was then tidally removed from its companion(s), leaving Cetus behind. Interactions during this scenario may plausibly explain Cetus’ SFH. Another consideration is that halo deformations due to the passage of massive satellites like the LMC and subsequent center of mass responses from massive host galaxies are not accounted for with the rigid potentials assumed in this work (see N. Garavito-Camargo et al., 2021; E. Patel et al., 2025). This phenomenon is beyond the scope of this analysis, but it is important to acknowledge that changing halo shape can alter orbital histories.